OpenStudy (anonymous):

Cameron opens a checking account and a savings account at his local bank. He deposits $10,000 into his checking account and $2,000 into his savings account. He will withdraw $1,200 from his checking account each year. His savings account earns 8% interest each year. In how many years will both accounts have the same balance? a. 5.5 years b. 5.6 years c. 5.7 years d. 5.8 years

OpenStudy (anonymous):

We replace the year's number with x 10000 - 1200x = 2000 + (2000*8/100)x -1200X -160x = -8000 -1360x = -8000 -1360x / 1360 = -8000/1360 - x = -5.88 x = 5.8 i wish that was helpful , i can explain more if required